# TELL-TALE: Task Efficient LLMs with Task Aware Layer Elimination

### v1

@@ Abstract
In this paper we introduce Tale, Task-Aware Layer Elimination, an inference-time algorithm that prunes entire transformer layers in an LLM by directly optimizing task-specific validation performance. We evaluate Tale on 9 tasks and 5 models, including LLaMA 3.1 8B, Qwen 2.5 7B, Qwen 2.5 0.5B, Mistral 7B, and Lucie 7B, under both zero-shot and few-shot settings. Unlike prior approaches, Tale requires no retraining and consistently improves accuracy while reducing computational cost across all benchmarks. Furthermore, applying Tale during finetuning leads to additional performance gains. Finally, Tale provides flexible user control over trade-offs between accuracy and efficiency. Mutual information analysis shows that certain layers act as bottlenecks, degrading task-relevant representations. Tale’s selective layer removal remedies this problem, producing smaller, faster, and more accurate models that are also faster to fine-tune while offering new insights into transformer interpretability.
@@ 1 Introduction
While Large Language Models (LLMs) have achieved remarkable success, they come with large computational overheads. This may prevent users in resource limited organizations or with high-throughput applications from using more capable models. This limitation has spawned an important research area on model compression. We focus here on the branch known as model pruning. Although it often reduces computational overhead, pruning typically requires complex implementation procedures, extensive retraining or fine-tuning, and may result in unpredictable performance degradation (Section 2 details related work).
In contrast, having observed that not all layers contribute to a particular task, we propose a lightweight, greedy, iterative layer pruning algorithm, Task Aware Layer Elimination (Tale). Tale operates at inference time, is hardware agnostic, directly optimizes for task-specific accuracy at each pruning step and consistently offers improved results over the original model. This improvement persists in interactions with fine tuning on our tasks. As illustrated in Figure 1 and detailed in Section 3, Tale systematically evaluates all possible single-layer removals at each iteration, selecting the layer whose elimination results in the highest validation accuracy. This process continues iteratively until performance improvements fall below a predefined threshold, ensuring that only layers with minimal or negative impact on task performance are removed.
Tale improves task specific accuracy and provides moderate computational reductions with minimal implementation complexity. This can be extremely useful for agentic systems with LLMs for dedicated tasks; TALE will make those LLMs perform better and more quickly. Our layer-wise pruning leverages the modular nature of transformer architectures, where each layer performs a complete transformation of the input representation through attention and feedforward mechanisms. This architectural property enables the removal of entire layers without requiring modifications to the remaining network structure.
We evaluate Tale with five LLMs, LLaMA 3.1 8B, Qwen 2.5 7B, Qwen 2.5 0.5B, Mistral 7B and Lucie 7B, on 9 diverse benchmark datasets (Sections 4 and 5) both in zero-shot and few-shot settings. We also investigated interactions between pruning with Tale finetuning. We find consistent improvements in both accuracy and computational efficiency in all cases. This suggests pre-trained LLMs contain redundant or even detrimental layers for a given downstream task. We analyze layer flow using the notion of mutual information (MI) to support this hypothesis (Section 6). Additionally, our experiments show Tale’s potential as a tool for understanding layer function in and across models (Section 7), thus aiding model interpretability.
@@ 2 Related Work
Zhu et al. (2024) distinguishes four primary approaches to model compression: model pruning, quantization, low-rank approximation, and knowledge distillation. Our work focuses on pruning, which can be categorized into unstructured, structured, and semi-structured methods. Unstructured pruning removes individual parameters, resulting in irregular, sparse structures Han et al. (2015b); Chen et al. (2015); Srinivas & Babu (2015); structured pruning eliminates entire components such as neurons, attention heads, or layers while maintaining the overall network structure He et al. (2017); Voita et al. (2019); Lagunas et al. (2021); Men et al. (2024). Semi-structured pruning combines fine-grained control with structural regularity, and has been explored in recent work Li et al. (2023); Frantar & Alistarh (2023b); Sun et al. (2024). Early pruning methods leveraged second-order information for structured pruning LeCun et al. (1989); Hassibi et al. (1993), but the field has since shifted toward computationally simpler, magnitude-based approaches that prune parameters by importance scores Han et al. (2015a); See et al. (2016); Narang et al. (2017)
Model pruning has benefited from information-theory. Building on the information bottleneck view of deep networks, Tishby et al. (2000); Tishby & Zaslavsky (2015) conceptualize layers as filters that preserve task-relevant information while discarding irrelevant input. This has led to model pruning proposals based on mutual information (MI). Fan et al. (2021) propose a layer-wise strategy that leverages MI estimates to reduce hidden dimensionality in a top-down manner. MI-guided approaches demonstrate that information-theoretic criteria can balance representation quality with model compression (Ganesh et al., 2020; Westphal et al., 2024). A central challenge, however, is the difficulty of estimating MI. Neural estimators such as MINE Ishmael Belghazi et al. (2018) provide principled solutions for high-dimensional settings, but in practice, probing classifiers Belinkov (2022) remain the dominant tool due to their efficiency and interpretability.
With regards to large transformers, Zhang & Papyan (2025) proposes a pruning strategy based on matrix approximations. Xia et al. (2023) demonstrates that pruning 7B models into 1.3B and 2.7B variants via structured layer and hidden-dimension pruning, combined with dynamic batch loading for fine-tuning, can yield submodels outperforming comparably sized models trained from scratch, though not matching the original model. Complementary work Frantar & Alistarh (2023a); Zhang et al. shows that contiguous blocks, especially attention layers, can often be pruned with limited performance loss, although critical “cornerstone” layers Zhang et al. remain essential for reasoning. Kim et al. (2024) investigates lock-level pruning based on weight importance, but these methods typically require fine-tuning techniques to recover baseline accuracy.
Most pruning methods compress models while attempting to match the original performance but often suffer accuracy degradation, which requires retraining or fine-tuning Xia et al. (2024). Improvements are generally measured relative to small models rather than original unpruned baselines. To the best of our knowledge, Tale is the only pruning method reported to improve over the original model’s performance without requiring any additional training or fine-tuning.
Figure 1: Illustration of Tale layer elimination. Candidate layers (yellow) are tested for removal, and the best-performing ones above the threshold are permanently dropped (red) until no further improvement is possible.
3 Tale , our Greedy-Selection Algorithm
3.1 Basics and Intuitions
@@
Our initial intuition for Tale came from examining the behavior of partial forward passes. Let $h^{(k)}$ denote the hidden representation after $k$ layers. Instead of always decoding from the final representation $h^{(\text{L})}$, we projected intermediate representations $h^{(k)}$ for $k<\text{L}$ directly into the vocabulary space using the output projection $W_{\text{out}}$, i.e.,
@@
We then compared the performance of $\hat{y}^{(k)}$ across different values of $k$. Surprisingly, we observed that for many tasks, intermediate layers ($k<\text{L}$) achieved higher accuracy than the final layer L (Figure 5). This indicated that additional depth does not always translate into better task-specific performance: some layers contribute marginally, while others introduce representational noise.
This experiment motivated our central hypothesis: not all layers in an LLM are equally useful, and selectively removing redundant layers can preserve—or even improve—downstream accuracy. Tale (Task-Aware Layer Elimination) formalizes this intuition into a principled, iterative pruning strategy.
3.2 Tale
Tale is a greedy iterative layer pruning algorithm for pre-trained open-weights LLM compression that systematically removes layers while preserving or even improving model performance (Algorithm 6). Starting with a full pre-trained model, Tale evaluates all possible single-layer removals at each iteration, computing the validation accuracy for each candidate pruned architecture. The layer whose removal results in the highest accuracy is permanently eliminated from the model, and this compressed architecture becomes the baseline for the next iteration. This process continues iteratively until the performance improvement falls below a predefined threshold, at which point the algorithm terminates and returns the most compressed model that maintains performance above the specified threshold. Our approach directly optimizes for task-specific accuracy at each pruning step, ensuring that only layers with minimal impact on the target objective are removed. This exhaustive evaluation strategy, while computationally intensive during the pruning phase, provides strong empirical guarantees about the optimality of each pruning decision within the greedy framework.
Algorithm 1 Tale : Greedy Iterative Layer Pruning
@@ 4 Benchmarks and Datasets
We evaluate Tale across a diverse suite of nine benchmarks spanning reasoning, language understanding, and commonsense knowledge. For mathematical reasoning, we include GSM8K-Hard, a curated subset of GSM8K Cobbe et al. (2021) with more than five premises per question to increase difficulty, and MATH500 Hendrycks et al. (2021b), a benchmark for symbolic and arithmetic reasoning (for evaluation details see Appendix A). For language understanding, we consider MMLU Hendrycks et al. (2021a) and BoolQ Clark et al. (2019), while Winogrande Sakaguchi et al. (2021), CommonsenseQA Talmor et al. (2019), and BIG-Bench Srivastava et al. (2023) capture commonsense and multi-task generalization. Finally, we include both ARC-Easy and ARC-Challenge Clark et al. (2018), which evaluate scientific and factual reasoning at varying difficulty levels. Together, these nine datasets cover a broad spectrum of downstream challenges and allow us to assess both the generality and task-specific benefits of our pruning strategy.

### v2

@@ Abstract
Large Language Models (LLMs) are typically deployed using a fixed architecture, despite growing evidence that not all layers contribute equally to every downstream task. In this work, we introduce TALE (Task-Aware Layer Elimination), an inference-time method that improves task performance by selectively removing layers that are irrelevant or detrimental for a given task. TALE optimizes task-specific validation performance, yielding a task-adapted architecture without retraining or modifying model weights. Across 9 tasks and 5 model families, under both zero-shot and few-shot settings, we show that TALE consistently matches or surpasses baseline performance while simultaneously reducing computational cost, outperforming general and layer-wise pruning approaches such as SLEB. Beyond inference-time gains, TALE synergizes with fine-tuning and few-shot learning, where task-adapted architectures lead to additional performance improvements. Computing TALE for a new task requires modest resources (1–2 GPU hours on an A100), making it a practical and deployable solution for task-specialized LLM inference.
TELL-TALE: Task Efficient LLMs with Task Aware Layer Elimination
Omar Naim1thanks: Equal contribution.   Krish Sharma [^1] footnotemark: 1   Niyar R Barman 1   Nicholas Asher  [^1] footnotemark: 1 1 IRIT, Toulouse , France
@@ 1 Introduction
The substantial computational costs of Large Language Models (LLMs) can prevent resource-constrained organizations and those with high-throughout applications from leveraging more capable models. This has led to a search for methods that boost task-specific performance and reduce computation costs. The use of multi-agent systems, with LLMs specialized for a particular role, has intensified the need for such methods. Yet finding such methods has proved elusive. Fine-tuning can increase task performance but does not reduce inference costs and requires significant training overhead and data. General pruning reduces computation costs but typically demands significant retraining and often results in substantial performance degradation on downstream tasks.
Tale, Task Aware Layer Elimination addresses this need. Tale increases task performance while modestly reducing computational overhead. Tale evaluates all possible single-layer removals at each iteration, selecting the layer whose elimination results in the highest validation accuracy. This process repeats until performance falls below a user predefined threshold.
Tale is a simple, hardware agnostic, greedy algorithm requiring no retraining that runs at inference time. Peer et al. (2022)’s similar greedy approach, on the other hand, requires retraining. Unlike all pruning methods of which we are aware, Tale optimizes for task-specific accuracy at each pruning step. We believe this is why Tale significantly improves results over the original model and interacts synergistically with fine-tuning.
Layer analysis on earlier models observed that not all layers contribute equally to a given taskDe Vries et al. (2020); Dalvi et al. (2020); Sajjad et al. (2023). Our study of state of the art, middle-sized transformers sharpens these observations: the layers most or least useful for a given LLM are highly dependent on the target task. These findings validate the attractiveness of Tale’s simple, task-specific, accuracy-driven design.
We showcase Tale ’s gains on five modern LLMs, LLaMA 3.1 8B, Qwen 2.5 7B, Qwen 2.5 0.5B, Mistral 7B and Lucie 7B, with 9 diverse benchmark datasets (Sections 4 and 5). We also compare Tale to previous training-free pruning methods on two older models (Llama 2 7B and 13B) and show that Tale achieves substantially higher accuracy with efficiency gains. With these benchmarks and model families, we show that removing task-misaligned layers can reduce inference cost and, in many cases, improve accuracy. Importantly, these improvements arise without retraining, architectural modification, or changes to the underlying model weights, and persist under alternative data splits, random seeds, and evaluation protocols. TALE complements fine-tuning and applies post hoc to existing checkpoints, making it a practical tool for deploying task-specialized LLMs.
Figure 1: How TALE works to produce BEST and Best Speedup above Baseline (BSBA) models.
@@ 2 Related work
While there is considerable work on model speedup through various pruning methods (pruning sparsity, acceleration, model compression), very few methods concentrate on boosting accuracy and almost no work has used compression methods to improve performance, which is what Tale does.
Structured pruning, which eliminates entire components such as neurons, attention heads, or layers He et al. (2017); Voita et al. (2019); Lagunas et al. (2021) is relevant to Tale . Frantar and Alistarh (2023a); Zhang et al. (2024) prune contiguous blocks with minimal performance loss. Similar to earlier methods that use representation similarity Dalvi et al. (2020), SLEB Song et al. (2024) removes entire layers based on the cosine similarity of their representations, which is the closest layer-wise baseline to Tale . SLEB’s layer selection is validated using a general metric (perplexity) on a linguistic dataset. SliceGPT Ashkboos et al. (2024) prunes layer dimensions via Principal Component Analysis. However, representational similarity, whether measured by cosine similarity, Center Kernel Alignment (Dalvi et al., 2020), or variance under PCA (Sajjad et al., 2023) is not a good measure for assessing a layer’s importance for a task. In Section 5.1 we show that representational similarity and validation with perplexity are sub-optimal for maximizing performance on specific downstream tasks.
SparseGPT Frantar and Alistarh (2023b) and Wanda Sun et al. (2023) perform weight-level unstructured pruning based on local reconstruction criteria or magnitude-activation products, achieving high sparsity but often degrading task-specific reasoning and linguistic abilities. Methods like OWL(Yin et al., 2024) and FLAP (An et al., 2024) use non uniform sparsity pruning to achieve important efficency gains, while Tang et al. (2025) uses an evolutionary algorithm and a fitness criterion on a selection of data to find models with a given sparsity. None of these approaches provide real improvements on accuracy for standard benchmarks, and for the most part they show significantly decreased performance from the baseline.
Peer et al. (2022) use a preset accuracy goal and a greedy algorithm like ours and show that their greedy-layer pruning (GLP) outperforms Top-layer pruning Dalvi et al. (2020) in almost all cases. Although superficially similar, GLP is very different from Tale. First, GLP retrains the pruned model on a downstream task T and compare the retrained model’s results on T to the those of the baseline model. Tale, on the other hand, involves no retraining. Even with retraining GLP’s reported accuracies for GLP models are far below our best models. Second, GLP restricts its algorithm to remove an a priori number of layers irrespective of optimization. Tale selects the number of layers that optimizes performance without setting a number a priori of layers to remove.
Tale differs from other training free methods in granularity and optimization objective: Tale is a structured, layer-level method that is explicitly optimized for task-specific accuracy, whereas the aforementioned training-free baselines are designed for compression using predefined limits and linguistic or reconstruction criteria. Section 5 compares various approaches with Tale , showing that Tale consistently and substantially improves accuracy over the unpruned baseline, a benefit not reliably achieved by other methods.
3 Basics and Intuitions
@@
Our intuition for Tale came from examining the behavior of partial forward passes. Let $h^{(k)}$ denote the hidden representation after $k$ layers. Instead of always decoding from the final representation $h^{(\text{L})}$, we projected intermediate representations $h^{(k)}$ for $k<\text{L}$ directly into the vocabulary space using the output projection $W_{\text{out}}$, i.e.,
@@
When comparing the performance of $\hat{y}^{(k)}$ across different values of $k$, we surprisingly observed that for many tasks, intermediate layers ($k<\text{L}$) achieved higher accuracy than the final layer L (Figure 7). Thus, additional layers did not always improve task-specific performance: some layers contribute marginally, while others introduce representational noise. Not all layers in an LLM are equally useful, and selectively removing redundant layers can preserve—or even improve—downstream accuracy. Tale (Task-Aware Layer Elimination) formalizes this intuition into a principled, iterative optimization strategy.
Figure 2: Accuracy progression of Tale across 3 benchmark datasets for LLaMA 3.1 8B. Each curve represents the accuracy at successive iterations. The $\star$ denotes the best-performing layer drop configuration, while the $\bullet$ highlights the Best Speed up with at least Baseline Accuracy (BSBA) configuration. Plots for all tasks are in Appendix E.
Algorithm 1 Tale : Iterative Layer Pruning
@@
3.1 Tale
Tale is a greedy, iterative layer pruning algorithm for pre-trained open-weights LLM compression that systematically removes layers while preserving or even improving model performance (Algorithm 1). Starting with a full pre-trained model, Tale evaluates all possible single-layer removals at each iteration, computing the validation accuracy for each candidate pruned architecture. The layer whose removal results in the highest accuracy is permanently eliminated from the model, and this compressed architecture becomes the baseline for the next iteration. This process continues iteratively until the performance improvement falls below a predefined threshold. For this paper we take baseline task accuracy - 8% to be the threshold, which allows the accuracy during our search to go slightly below baseline to capture eventual, substantial increases. We have found no cases where the trajectory later goes above the baseline. Once the threshold is reached, the algorithm terminates and returns the most compressed model with performance above the threshold.
Task performance optimization trajectories Figure 2 visualizes three iterative model optimization processes for LLaMA 3.1 8B with Tale . Each curve tracks accuracy as layers are progressively removed. The graphs reveal a general pattern across almost all our tasks: the first iteration of TALE typically provides a large boost in accuracy, followed by slight increases or decreases; they then follow a monotonic decreasing path to accuracies below the baseline and eventually to 0. A full set of curves is in Figure 11. The curves in themselves are worthy of future study.
Three consistent patterns emerge from these trajectories: (i) Tale identifies compressed BEST models that outperform the original across diverse tasks, with $\star$ markers lying strictly above baseline. Best Speed up with at least Baseline Accuracy (BSBA) are the models with best speed up above Tale’s stopping point (with $\bullet$). (ii) Accuracy improvements persist across multiple pruning steps before diminishing returns, showing that substantial redundancy exists even in carefully tuned pretrained models. (iii) Pruning dynamics are task-specific: datasets such as ARC-Easy and MMLU tolerate deeper pruning while continuing to improve, whereas reasoning-heavy tasks like GSM8K-Hard converge earlier, reflecting heterogeneous layer importance across domains.
@@ 4 Benchmarks and Datasets
We evaluate Tale across a diverse suite of nine benchmarks spanning reasoning, language understanding, and commonsense knowledge. For mathematical reasoning, we use GSM8K-Hard, a curated subset of GSM8K Cobbe et al. (2021) with more than five premises per question to increase difficulty, and MATH500 Hendrycks et al. (2021b). (for evaluation details see Appendix A). For language understanding commonsense reasoning and multi-task generalization, we use MMLU Hendrycks et al. (2021a) and BoolQ Clark et al. (2019); Winogrande Sakaguchi et al. (2021), CommonsenseQA Talmor et al. (2019), andBIG-Bench Srivastava et al. (2023), ARC-Easy and ARC-Challenge Clark et al. (2018).
In our study Tale requires only modestly-sized sets for task-specific optimization, ranging from 500 to 1500 examples. As seen in Table 3 (Appendix C), once the validation set size exceeds 500 examples, the set of layers dropped stabilizes across all tasks. For each benchmark, we partition the available data into three disjoint subsets: (i) a training pool, (ii) a held-out optimization split drawn from the training pool, and (iii) a final evaluation split (the official test set when available, or a held-out validation set otherwise).
TALE uses only the optimization split to select which layers to remove for a given task. The evaluation split is never used during layer selection. All results reported throughout the paper correspond to performance on this unseen evaluation split.
When official test sets are unavailable, we follow prior work and report results on held-out validation splits that are disjoint from the data used for layer optimization. To see results with two different training subsets see Tables 2 and 17.